### v1

# Normality in the square of the Sorgenfrey line

Abstract.
We consider sets of reals $X$ endowed with the Sorgenfrey lower limit topology denoted $X[\leq]$. Przymusiński proved that if $X$ is a $Q$-set then $(X[\leq])^{2}$ is normal. While the converse is not in general true we consider examples of sets of the reals for which $(X[\leq])^{2}$ is normal or just pseudo-normal. For example, if $X$ is a $\lambda$ set, then $(X[\leq])^{2}$ is pseudo-normal but assuming CH there is an $X$ concentrated on a countable dense subset (so not a $\lambda$-set) but still $(X[\leq])^{2}$ is normal.
Key words and phrases:
$Q$-set, $\lambda$-set, Sorgenfrey topology
2020 Mathematics Subject Classification:
Primary:
1. Introduction
A $Q$-set is a subset of the reals with the property that every subset is $G_{\delta}$ (equivalently $F_{\sigma}$) in the relative subspace topology. The notion was first introduced in [hausdorff1933] and $Q$-sets were used to produce consistent counterexamples to the Normal Moore Space Conjecture [Bing1951]. For example, if $X\subseteq 2^{\omega}$, the set of branches $b_{x}=\{x\upharpoonright n\}$ in $2^{<\omega}$ is an almost disjoint family and the corresponding $\Psi$-space on $2^{<\omega}$ is normal if and only if $X$ is a $Q$-set. Similarly, if $X$ is a subset of the $x$-axis in the upper half plan of ${\mathbb{R}}^{2}$, then the Moore-Niemitzky Plane topology (aka the bubble space) is normal if and only if $X$ is a Q-set. For more on these examples see [T]. On the other hand Prymusiński [T.Przymusinski1973] proved that if $X$ is a Q-set endowed with the Sorgenfrey Topology then $X^{2}$ is normal.
Recall, the Sorgenfrey line is the real line endowed with the topology, also known as the lower limit topology, generated by using the half-open intervals $[a,b)$ as a basis. For a set of reals $X$ we will denote $X[\leq]$ as the resulting topology taking $X$ as a subspace of the Sorgenfrey line.
We first observe that the converse to Przymusiński’s theorem does not hold: Todorčević observed that if $X$ is an entangled set of reals and is endowed with the Sorgenfrey topology, then every finite power is both hereditarily Lindelöf and hereditarily separable [Todorcevic1989]. Recall that CH implies both the existence of entangled sets of reals and the non-existence of $Q$-set.
There is the weaker notion of a set of reals $X$ being a $\lambda$-set. $X$ is a $\lambda$-set if every countable subset is $G_{\delta}$. While the existence of a $Q$-set of reals is independent, there are $\lambda$-sets in ZFC. Analagous to the $Q$-set results mentioned above, the $\Psi$-space over the set of branches of a subset $X\subseteq 2^{\omega}$ is pseudo-normal (meaning that every pair of closed sets, one of which is countable, can be separated by disjoint open sets) if and only if $X$ is a $\lambda$-set and also $X$ is a $\lambda$-set if and only if the Moore-Niemytzki plane over $X$ is pseudo-normal (see [vD])
We provide some partial converses to Przymusiński’s theorem that underscore what is needed to prove establish normality of $(X[\leq])^{2}$. We show that the expected analogous result for the Sorgenfrey topology on a $\lambda$-set holds. Namely, if $X$ is a $\lambda$-set of reals, then $(X[\leq])^{2}$ is pseudonormal. We also construct, assuming CH, a set of reals for which $X([\leq])^{2}$ is normal such that $X$ is concentrated on a countable dense subset, and so not a $\lambda$-set.
We will use the following terminology and notation throughout.
Definition 1.1.
Let $X$ be a topological space. Let $A,B$ be subsets in $X$, AA has the covering property with respect to $B$ iff there exists a countable cover $\{U_{n}\}_{n<\omega}$ such that $\overline{U_{n}}\cap B=\emptyset$. If there is no ambiguity in the context, we just say $A$ has the covering property.
Definition 1.2.
Let $z=(x,y)\in\mathbb{R}^{2}$, $r>0$ in $\mathbb{R}$, denote the basic Sorgenfrey neighbourhood of radius $r$ around $z$ to be $U(z,r)=[x,x+r)\times[y,y+r)$.
Definition 1.3.
Let $A,B\subseteq\mathbb{R}$, $f:A\to B$, denote the graph of $f$ in $\mathbb{R}^{2}$ to be $G_{f}=\{(x,f(x)):x\in A\}$.
All of our topological terminology is standard and we refer the reader to Engelking for any undefined notions.
2. Q-sets and the Sorgenfrey square
We begin by recalling a simple fact about monotone functions that we will use to establish a partial converse to Przymusiński’s theorem.
Lemma 2.1.
Let $A,B$ be subsets of reals. For any strictly decreasing function $f:A\to B$, there is a countable $A_{0}$ such that $f\upharpoonright(A\setminus A_{0}):A\setminus A_{0}\rightarrow f(A\setminus A_{0})$ is a homeomorphism.
Lemma 2.2.
Suppose $X$ is a set of reals. Suppose further that
  1. (1)
    there exists a strictly decreasing homeomorphism $f:X\to X$, and
  2. (2)
    letting $A\subseteq X$, define
    $B$ $=X\setminus A$
    $G_{f|_{A}}$ $=\{(a,f(a)):a\in A\}$
    $G_{f|_{B}}$ $=\{(b,f(b):b\in B)\}.$
If there exists $\mathcal{U},\mathcal{V}\subseteq X^{2}$ Sorgenfrey open such that $\mathcal{U}\supseteq G_{f|_{A}},\mathcal{V}\supseteq G_{f_{B}}$ and $\mathcal{U}\cap\mathcal{V}=\emptyset$, then $A$ is relatively $F_{\sigma}$ in $X$.
Proof.
By assumption, let $\mathcal{U},\mathcal{V}$ be Sorgenfrey open sets such that $\mathcal{U}\supseteq G_{f|_{A}},\mathcal{V}\supseteq G_{f|_{B}}$.
Note that $A=\cup_{n<\omega}A_{n}$ where $A_{n}=\{a\in A:U((a,f(a)),1/n)\subseteq\mathcal{U}\}$, we claim that
Claim 2.3.
$\overline{A_{n}}^{E}\cap X\subseteq A$
Proof.
Assume otherwise, if $a_{n}\to b$ for $a_{n}\in A$ and $b\in B$. There exists an $r>0$ such that $U((b,f(b)),r)\subseteq\mathcal{V}$. By continuity of $f$, we have $f(a_{n})\to f(b)$ as well. Let $\epsilon=\min(1/n,r)$, Wlog there exists $a_{n_{0}}$ such that $b<a_{n_{0}}<b+\epsilon$ and $f(b)<f(a_{n_{0}})<f(b)+\epsilon$. Since $\epsilon<1/n$, $(a_{n_{0}},f(a_{n_{0}})+\epsilon)$ is a point in $U((a_{n_{0}},f(a_{n_{0}})),1/n)\subseteq\mathcal{U}$ However, we have
$f(b)<f(a_{n_{0}})+\epsilon<f(b)+2\epsilon<f(b)+r$
So $(a_{n_{0}},f(a_{n_{0}}))$ is also a point in $U((b,f(b)),r)\subseteq\mathcal{V}$, but $\mathcal{U}$ and $\mathcal{V}$ should have empty intersection hence this is a contradiction. ∎
Now with this claim, we have $A=\cup_{n<\omega}(\overline{A_{n}}^{E}\cap X)$ which is a relatively $F_{\sigma}$ set. ∎
We now prove the following partial converse to Przymusiński’s theorem:
Proposition 2.4.
Let $(X[\leq])^{2}$ be normal. If there exists a strictly decreasing homeomorphism $f:X\to X$, then $X$ is a $Q$-set.
Proof.
Since $f$ is a homeomorphism, its graph $G_{f}$ is closed and discrete. Hence any two subsets of $G_{f}$ is closed. In particular, take any $A\subseteq X$, $G_{f|_{A}}$ and $G_{f|_{X\setminus A}}$ can be separated by open sets. By 2.2, $A$ is relatively $F_{\sigma}$ in $X$. Hence $X$ is a $Q$-set. ∎
The following is a slightly more general partial result:
Proposition 2.5.
Let $(X[\leq])^{2}$ be normal. If there exists a strictly decreasing function $f:X\to X$ (not necessarily continuous with respect to the Euclidean topology), then $X\setminus C$ is a $Q$-set for some countable subset $C$.
Proof.
By 2.1, there exists a countable set $C$ such that $f|_{X\setminus C}:X\setminus C\to X\setminus C$ is a homeomorphism. Then let $g=f|_{X\setminus C}$, the graph $G_{g}$ of $g$ is closed and discrete. Since $(X[\leq])^{2}$ is normal, any two subsets of $G_{g}$ can be separated by Sorgenfrey open subsets of $X^{2}$. In particular, let $X^{\prime}=X\setminus C$, take any subset $A\subseteq X^{\prime}$, $G_{g|_{A}}$ and $G_{g|_{X\prime\setminus A}}$ can be separated by open sets $\mathcal{U},\mathcal{V}$. Take their intersection to $(X^{\prime})^{2}$, $\mathcal{U}\cup(X^{\prime})^{2}$ and $\mathcal{V}\cup(X^{\prime})^{2}$ are still Sorgenfrey open in $(X^{\prime})^{2}$. By 2.2, $A$ is relatively $F_{\sigma}$ in $X^{\prime}$, hence $X^{\prime}$ is a $Q$ set, and it only lacks countably many points. ∎
Baumgartner proved that PFA implies that any two $\aleph_{1}$-dense sets of reals are order isomorphic [Baumgartner1973]. This consequence of PFA has come to be referred to as Baumgartner’s Axiom (BA). From the proof of the above proposition we obtain the following corollary:
Corollary 2.1.
BA implies that if $X$ is an uncountable set of reals and $(X[\leq])^{2}$ normal then $X$ contains an uncountable $Q$-set.
Proof.
Any uncountable $X$ contains an $\aleph_{1}$-dense subset. By BA it is order isomorphic to its reverse order. So by the proof of Proposition 2.5 there is a countable $C$ such that $X\setminus C$ must be a $Q$-set.∎
For the remainder of this section we focus on constructing, from CH, an uncountable set of reals whose square, under the Sorgenfrey topology, is normal. Of course, under CH, no uncountable set of reals is a Q-set and so the converse of Przymusiński’s theorem fails assuming CH.
We begin by recalling another simple folklore result on monotone real-valued functions. For completeness sake, we provide a proof.
Lemma 2.6.
Let $f:A\to\mathbb{R}$ be strictly decreasing on uncountable $A\subseteq\mathbb{R}$, then $f$ can be extended to strictly decreasing function on $\overline{A}\setminus C$ for some countable set $C$.
Proof.
$f$ can be extended to a decreasing function on $\overline{A}$. Note that this function is not necessarily strictly decreasing. By 2.1, we can remove a countable set $B_{0}$ such that $f$ is homeomorphism on $\overline{A}\setminus f^{-1}(B_{0})$ hence strictly decreasing. We will show $f^{-1}(B_{0})$ is countable by showing for every $b\in B_{0}$, $f^{-1}(b)$ has at most 3 points. Assume there are 4 points $x<y<z<w$ in $f^{-1}(b)$. Note that between $x$ and $w$ there is at most 1 point in $A$, since for every $z_{1}<z_{2}\in A$,
$f(x)\geq f(z_{1})\geq f(z_{2})\geq f(w)=f(x)$
but $f$ is strictly decreasing on $A$, so there can only be at most 1 point in $A$. Hence, there are no points in $\overline{A}\setminus A$. But $y,z$ can not be both in $A$, we have a contradiction. Let $C=f^{-1}(B_{0})$, we have the result. ∎
Recall a set $L\subseteq\mathbb{R}$ is $\mathbf{n}$-entangled iff for any set of disjoint $n$-tuples $A\subseteq L^{n}$ of size $\omega_{1}$ and for every $s\in 2^{n}$ there exists $x,y\in A$ such that $\forall i\leq n(x_{i}<y_{i})\iff s_{i}=0$. $L$ is entangled iff $L$ is $\mathbf{n}$-entangled for all $n<\omega$. Let $s\in 2^{n}$, $A\subseteq L^{n}$ is said to have type $s$ iff $\forall i\leq n(x_{i}<y_{i})\iff s_{i}=0$.
The following lemma is due to [Todorcevic1985].
Lemma 2.7.
[Todorcevic1985] Let $E\subseteq\mathbb{R}$ to have the following property:
  • No one-to-one continuous function can be defined from $A\to E$ for any uncountable $A\subseteq E^{n}$, for any $n<\omega$
Then $E$ is entangled.
Proof.
Assume $E$ is not entangled, let $m<\omega$ be the least such that $E$ is $m$-entangled, but not $m+1$-entangled. Then there exists an uncountable $A\subseteq E^{m+1}$ such that $A$ is missing one type $s\in 2^{n}$, with $s=(s_{1},s_{2},\dots,s_{m},s_{m+1})$. Wlog, let $s_{m+1}=0$. Define $f:E^{m}\to E$ by $f(z)=r$ iff $z\text{\textasciicircum}r\in A$. We will show $f$ is continuous on some uncountable subset of $dom(f)$, which is not possible by the property $E$ has. For any $p\in dom(f)$, define
$\omega(p)=\cap\{\overline{f(dom(f)\cap I)}:I\text{ open in }E^{m},p\in I\}$
$\omega(p)$ is all the possible limit at the point $p$, if $|\omega(p)|=0$ or $1$, $f$ is continuous at $p$. Let $B=\{p\in dom(f):|\omega(p)|\geq 2\}$, assume $B$ is uncountable. By countable chain condition and symmetry, there exists an uncountable subset $B_{0}$ such that
$\forall z\in B_{0},\exists r\in\omega(z)\;f(z)<d<r$
Note that $B_{0}\subseteq E^{m}$, and since $E$ is $m$-entangled, there exists $u,v\in B_{0}$, $u=(u_{1},\dots,u_{m}),v=(v_{1},\dots,v_{m})$, such that $\forall i\leq m(u_{i}<v_{i})\iff s_{i}=0$. Let
$I=\{z\in E^{m}:\forall i\leq m(u_{i}<z_{i})\iff s_{i}=0\}$
II is an open set. However for all $z\in dom(f)\cap I$, since $s_{m+1}=0$ and $A$ is missing this type, we can only have $f(z)<f(u)<d$, and $\omega(v)\subseteq\overline{f(dom(f)\cap I)}\implies\omega(v)<d$ which is a contradiction. ∎
Theorem 2.8.
Under CH, there exists a set $E$ such that
  1. (1)
    No one-to-one continuous function can be defined from $A\to E$ for any uncountable $A\subseteq E^{n}$, for any $n<\omega$.
  2. (2)
    $E$ contains $\mathbb{Q}$ and $\mathbb{Q}$ is not a relative $G_{\delta}$-set in $E$. Moreover, $E$ can be chosen so that it is concentrated on $\mathbb{Q}$. I.e., $E\setminus U$ is countable for every open set $U\supseteq\mathbb{Q}$.
Proof.
Define
$\mathcal{F}=\{$ $f:U\to\mathbb{R}:f\text{ one-to-one continuous function on Borel set }U\subseteq\mathbb{R}^{n},n<\omega,$
$\text{and let }U=\Pi_{i=1}^{n}U_{i},\text{we have }\cup_{i=1}^{n}U_{i}\text{ uncountable}\}$
$\mathcal{A}=\{$ $A\subseteq\mathbb{R}:A\text{ is open},\ A\supseteq\mathbb{Q}\}$
Note that both $\mathcal{F},\mathcal{A}$ have cardinality $\mathfrak{c}$, so under $CH$ we can enumerate them in $\omega_{1}$ as follows
$\mathcal{F}=\{f_{\alpha}:\alpha<\omega_{1}\},\mathcal{A}=\{A_{\alpha}:\alpha<\omega_{1}\}.$
For every $\alpha<\omega_{1}$ define the following set:
$C_{\alpha}$ $=\bigcap_{\beta\leq\alpha}A_{\alpha}\setminus\mathbb{Q}.$
Note that $C_{\alpha}$ is uncountable by the Baire Category Theorem. We now define elements $x_{\alpha}\in\mathbb{R}$ recursively as follows:
Stage 0: Take $x_{0}\in C_{0}$ such that $f_{0}(\{x_{0}\}^{n})\not=x_{0}$
Stage $\bm{\alpha<\omega_{1}}$: We have $x_{\beta}$ defined already for all $\beta<\alpha$, define
$X_{\alpha}$ $=\{x_{\beta}:\beta<\alpha\}$
$I_{\alpha}$ $=\{f_{\beta}(x_{\gamma_{1}},x_{\gamma_{2}},\dots,x_{\gamma_{n}}):\beta\leq\alpha,\gamma_{i}<\alpha,i\leq n\}$
$P_{\alpha}$ $=\cup\{f_{\beta}^{-1}(x_{\gamma}):\beta\leq\alpha,\gamma<\alpha\}$
Note that all of the above sets are countable, and $C_{\alpha}$ is uncountable. Hence, we can choose $x_{\alpha}\in C_{\alpha}\setminus(X_{\alpha}\cup I_{\alpha}\cup P_{\alpha})$, such that either $x_{\alpha}\not\in\cup dom(f_{\alpha})$ or $f_{\alpha}([X_{\alpha}]^{n}\cap dom(f))\not\ni x_{\alpha}$.
Finally we define $E=\mathbb{Q}\cup\{x_{\alpha}:\alpha<\omega_{1}\}$.
Verify Property 1: Fix $n<\omega$, for any one to one continuous function $g:A\to E$ defined on some uncountable set $A\subseteq E^{n}$, $g$ can always be extended to a continuous function on a Borel set. , i.e., there is an $\alpha_{0}$ such that $f_{\alpha_{0}}\upharpoonright A=g$. However, by choice of $x_{\alpha}$, for all $\alpha>\alpha_{0}$, we have that $f_{\alpha_{0}}(\{x_{\gamma}:\gamma<\alpha\}^{n})$ can not take on any values in $\{x_{\beta}:\beta>\alpha_{0}\}$. Indeed for all $\beta>\alpha_{0}$, if $\alpha_{0}<\beta<\alpha$, we have $x_{\alpha}\not\in P_{\alpha}\implies x_{\alpha}\not\in\cup f^{-1}(x_{\beta})\implies f_{\alpha_{0}}(\{x_{\gamma}:\gamma<\alpha\}^{n})\neq x_{\beta}$. If $\alpha_{0}<\alpha<\beta$, we have $x_{\beta}\not\in I_{\beta}\implies x_{\beta}\not\in f(\{x_{\gamma}:\gamma<\beta\}^{n})\implies x_{\beta}\not\in f(\{x_{\gamma}:\gamma<\alpha\}^{n})$. So for all $z\in E^{n}$ such that $z_{i}=x_{\alpha}$ for some $i\leq n$, $f_{\alpha_{0}}(z)\in\mathbb{Q}\cup\{x_{\beta}:\beta\leq\alpha_{0}\}$ which is a countable set. Hence $g^{-1}(E)$ is countable contradicting the fact that $g$ was defined on the uncountable set $A$.
Verify Property 2: That $X$ is concentrated on $\mathbb{Q}$ follows easily from the fact that for every $\alpha$, $X\setminus A_{\alpha}\subseteq\{x_{\beta}:\beta<\alpha\}$. In particular, $\mathbb{Q}$ is not $G_{\delta}$ in $X$ and so $X$ is not a $\lambda$-set.
By 2.7, the set $E$ we constructed in 2.8 is entangled and in particular $2$-entangled. By 2.6 this implies no monotone functions can be defined on $E$. Hence we have the following result.
Theorem 2.9.
Let $E$ be any subset of $\mathbb{R}$} that admits no uncountable strictly monotone functions (e.g., the set constructed in Theorem 2.8), then $(E[\leq])^{2}$ is normal.
Proof.
Let $A,B$ be closed in $(E[\leq])^{2}$. Recall the shoelace lemma: to separate $A$ and $B$ it suffices to cover each set by a countable family of open sets whose closures are disjoint from the other closed set. By symmetry, it suffices to find a countable cover $\{U_{n}\}_{n<\omega}$ of $A$ such that $\overline{U_{n}}\cap B=\emptyset$.
Recall we defined $A^{\prime}\subseteq A$ to have the covering property (with respect to $B$) if there is a countable open cover $\{U_{n}\}$ such that $\overline{U_{n}}\cap B=\emptyset$
Clearly, we have:
Lemma 2.10.
If $A^{\prime}\subseteq A$ is Lindelof, then $A^{\prime}$ has the covering property.
For all $z\in B^{c}$ choose $n_{z}\in\mathbb{N}$ such that $U(z,1/n_{z})\cap B=\emptyset$, this is possible since $B$ is Sorgenfrey closed. By letting $A_{1}=A\cap\overline{B}^{E}$ and $A_{2}=A\setminus\overline{B}^{E}$, we have
Observation 1.
$A=A_{1}\cup A_{2}$
Claim 2.11.
$A_{2}$ has the covering property.
Proof.
Note that $A_{2}\subseteq E\setminus cl_{E}(B)=U$ and the latter is Euclidean open. Since $E^{2}$ is hereditarily Lindelöf with respect to the Euclidean topology, $U$ can be written as the countable union of basic open sets $B_{n}$ such that $\overline{B_{n}}\cap B=\emptyset$ for each $n$ so, since the Sorgenfrey topology is finer than the Euclidean, $A_{2}$ has the covering property. ∎
Claim 2.12.
$A_{1}$ has the covering property.
Proof.
We will break down $A_{1}$ into countably many pieces, each of which is the graph of a monotone decreasing function and prove every piece has the covering property. Define
$A_{q}=\{z=(x,y)\in A_{1}:x+y<q<x+y+1/n_{z}\}$
Denote $l_{q}=\{(x,y)\in X^{2}:x+y=q\}$, $A_{q}$ is just the collection of points $z$ which has the straight line $l_{q}$ passing through the interior of its nbhd $U(z,1/n_{z})$. We see that
Observation 2.
$A_{1}=\cup_{q\in\mathbb{Q}}A_{q}$
So now it suffices to show that each $A_{q}$ has the covering property. First note that
Claim 2.13.
(non-increasing property) For all $z_{1}=(x_{1},y_{1}),z_{2}=(x_{2},y_{2})\in A_{q}$, if $x_{1}<x_{2}$, then $y_{2}\leq y_{1}$.
Proof.
For otherwise we have $x_{1}<x_{2}$ and $y_{1}<y_{2}$. Since the line $l_{q}$ passes through both nbhds $U(z_{1},1/n_{z_{1}}),U(z_{2},1/n_{z_{2}})$, this means $z_{2}$ is in the Euclidean interior of $U(z_{1},1/n_{z_{1}})$. However, $\operatorname{int}_{E}U(z_{1},1/n_{z_{1}})$ is a Euclidean open set, and since $z_{2}\in\overline{B}^{E}$, we have that $\operatorname{int}_{E}U(z_{1},1/n_{z_{1}})$ also contains a point from $B$, which is a contradiction, because we chose $n_{z_{1}}$ so that its nbhd $U(z_{1},1/n_{z_{1}})$ is disjoint from $B$. ∎
Claim 2.14.
The sets $M_{1}=\{x\in p_{0}(A_{q}):|p_{0}^{-1}(x)|>1\},M_{2}=\{x\in p_{1}(A_{q}):|p_{1}^{-1}(x)|>1\}$ are countable.
Proof.
For each $x\in M_{1}$, fix $\alpha_{x}<\beta_{x}\in p_{0}^{-1}(x)$. And for any $x_{1}<x_{2}\in M_{1}$, we have $\beta_{x_{2}}\leq\alpha_{x_{1}}$ by the previous claim. And so the intervals $(\alpha_{x_{1}},\beta_{x_{1}})$ and $(\alpha_{x_{2}},\beta_{x_{2}})$ are disjoint. Hence, we have an uncountable set of pairwise disjoint open intervals on $\mathbb{R}$}, contradiction. The proof for $M_{2}$ is the same. ∎
Let $L_{1}=p_{0}^{-1}(M_{1}),L_{2}=p_{1}^{-1}(M_{2})$, and $L=L_{1}\cup L_{2}$, $K=A_{q}\setminus L$, we see that
Observation 3.
$A_{q}=K\cup L$
Claim 2.15.
$L$ has the covering property.
Proof.
We know $L_{1},L_{2}$ are Lindelof since by Claim 2.14 each is a countable union of Lindelof subspaces. And by 2.10, $L$ has the covering property. ∎
Now it suffices to show $K$ has the covering property.
Claim 2.16.
$K$ has the covering property
Proof.
Note that $K$ is now a graph of some strictly decreasing function $g$. By the definition of $E$, such $g$ can only be defined on a countable set, hence the graph is countable hence Lindelof. ∎
This shows that each $A_{q}$ has the covering property completing the proof that $A_{1}$ has the covering property. ∎
Now with all the observations and claims, we obtain a countable cover of $A$ witnessing that $A$ has the covering property and so by the shoelace lemma, $A$ and $B$ can be separated by open sets, hence $E^{2}$ is normal. ∎
Now one may try to categorize all sets $E\subseteq\mathbb{R}$ such that $(E[\leq])^{2}$ into either $Q$-set or a $2$-entangled set. However, the following example shows that is not the case.
Theorem 2.17.
Assuming CH, there exists a set $E$ such that
  1. (1)
    $E$ is not 2-entangled.
  2. (2)
    $E$ is concentrated on $\mathbb{Q}$ (hence not a Q-set nor a $\lambda$-set).
  3. (3)
    $(E[\leq])^{2}$ is normal.
Proof.
Let $f:\mathbb{R}\to\mathbb{R}$ be any one-to-one increasing function such that $dom(f)$ is of cardinality $\aleph_{1}$ and $\{x\in dom(f):f(x)\neq x\}$ is of cardinality $\aleph_{1}$. Wlog, we can assume $f(x)\neq x$ for all $x\in dom(f)$. We will preserve this function $f$ while making it concentrated on rationals and killing all the strictly decreasing functions on uncountable sets. Let $G_{f}={(u_{\alpha},v_{\alpha}):\alpha<\omega_{1}}$. As in 2.8, let $\mathcal{F}=\{f_{\alpha}:\alpha<\omega_{1}\}$ be the set of strictly decreasing functions, $\mathcal{A}=\{A_{\alpha}:\alpha<\omega_{1}\}$ be the set of open sets containning $\mathbb{Q}$ and for each $\alpha$, $C_{\alpha}=\cap_{\beta\leq\alpha}A_{\alpha}\setminus\mathbb{Q}$. We now define elements $x_{\alpha}\in\mathbb{R}$ recursively as follows:
Stage 0: Take $x_{0}\in C_{0}\setminus\{u_{0},v_{0}\}$ such that $f_{0}(x_{0})\not=x_{0}$
Stage $\bm{\alpha<\omega_{1}}$: We have $x_{\beta}$ defined already for all $\beta<\alpha$, define
$X_{\alpha}$ $=\{x_{\beta}:\beta<\alpha\}\cup\{u_{\beta}:\beta<\alpha\}\cup\{v_{\beta}:\beta<\alpha\}$
$I_{\alpha}$ $=\{f_{\beta}(x):\beta\leq\alpha,x\in X_{\alpha}\}$
$P_{\alpha}$ $=\{f_{\beta}^{-1}(x):\beta\leq\alpha,x\in X_{\alpha}\}$
Note that all of the above sets are countable, and $C_{\alpha}$ is uncountable. Hence, we can choose $x_{\alpha}\in C_{\alpha}\setminus(X_{\alpha}\cup I_{\alpha}\cup P_{\alpha})$, such that either $x_{\alpha}\not\in dom(f_{\alpha})$ or $f_{\alpha}(X_{\alpha})\not\ni x_{\alpha}$.
Finally we define $E=\mathbb{Q}\cup\{x_{\alpha}:\alpha<\omega_{1}\}\cup\{u_{\alpha}:\alpha<\omega_{1}\}\cup\{v_{\alpha}:\alpha<\omega_{1}\}$.
Now $E$ is not entangled since the uncountable one-to-one increasing function $f$ exists. $(2)$ and $(3)$ follows exactly as in 2.8. ∎
3. $\lambda$-sets and the Sorgenfrey square
Theorem 3.1.
If $X$ is a $\lambda$-set, then $(X[\leq])^{2}$ is pseudonormal.
Proof.
Let $A,B$ be closed sets and assume $B$ is countable. We will again use shoelace lemma, and so suffices to show that $A$ has the covering property with respect to $B$. Since $B$ is countable it clearly has the covering property with respect to $A$. We begin by defining the following sets:
$B_{0}$ $=p_{0}(B)$
$B_{1}$ $=p_{1}(B)$
$A_{0}$ $=A\cap p_{0}^{-1}(B_{0})$
$A_{1}$ $=A\cap p_{1}^{-1}(B_{1})$
$W$ $=A\setminus(A_{0}\cup A_{1})$
Note that $B_{0},B_{1}$ are countable, so $A_{0}$ and $A_{1}$ are Lindelöf. Hence $A_{0}$ and $A_{1}$ have the covering property. Now it suffices to prove
Claim 3.2.
$W$ has the covering property
Proof of Claim.
We know that there are families $\{F_{n}:n\in\omega\}$ and $\{H_{n}:n\in\omega\}$ of relatively closed (in the Eucliden topology) subsets of $X$ such that
$p_{0}(W)\subseteq p_{0}(B)^{c}$ $=\cup_{n\in\omega}F_{n}$
$p_{1}(W)\subseteq p_{1}(B)^{c}$ $=\cup_{m\in\omega}H_{m}$
For any $z\in W$, there exists $s_{z}>0$ such that $U(z,s_{z})\cap B=\emptyset$. Define
$W_{n,m}=\bigcup\{U(z,s_{z}):\ z=(x,y)\in W,x\in F_{n},y\in H_{m}\}.$
We see $W\subseteq\cup_{n,m}W_{n,m}$ and each $W_{n,m}$ is Sorgenfrey open. We know that $W_{n,m}\cap B=\emptyset$, so it suffices to show $(\overline{W_{n,m}}\setminus W_{n,m})\cap B=\emptyset$.
Assume otherwise. Then there exists a $z\in B\setminus W_{n,m}$ such that for all $k\in\omega$, there exists a $w_{k}=(x_{k},y_{k})\in W_{n,m}$ such that, $U(z,1/k)\cap U(w_{k},s_{w_{k}})\neq\emptyset$. We have the following three cases depicted below.
  1. (1)
    For infinitely many $k$ we have $x_{k}<a,y_{k}<b$
    In this case, we have $(a,b)\in U(w_{k},s_{w_{k}})\subseteq W_{n,m}$, which is a contradiction since we chose $z=(a,b)$ to be disjoint from $W_{n,m}$
  2. (2)
    For infinitely many $k$ $x_{k}\to a$. But since each $x_{k}$ is in $F_{n}$, it follows that $a\in F_{n}$ which is impossible.
  3. (3)
    For infinitely many $k$ $y_{k}\to b$. But since $y_{k}$ is in $H_{m}$ it would follow that $b\in H_{m}$, also impossible.
Hence $\overline{W_{n,m}}\cap B=\emptyset$, we have the desired covering. ∎
From the previous example, we know that it is possible for $(X[\leq])^{2}$ to be normal, while $X$ is not a $\lambda$-set. However, we do have the following partial converse analogous to Proposition 2.4.
Proposition 3.3.
Let $(X[\leq])^{2}$ be pseudo-normal. If there exists a strictly decreasing homeomorphism $f:X\to X$ with respect to the Euclidean topology, then $X$ is a $\lambda$-set.
Proof.
Since $f$ is a homeomorphism, the graph $G_{f}$ is closed and discrete, any subset of $G_{f}$ is closed in $X^{2}$. Let $A$ be any countable subset, we have $G_{f|_{A}}$ is countable as well, hence $G_{f|_{A}}$ and $G_{f|_{X\setminus A}}$ are closed and can be separated by open sets. Then by 2.2, $X\setminus A$ is relatively $F_{\sigma}$ in $X$, hence $X$ is a $\lambda$-set. ∎
Finally, we close with some remarks and a question. As we have seen, the converse to Przymusiński’s theorem is independent (e.g., assuming CH). Is the converse consistent? I.e., is it consistent that for an uncountable subset of the reals $X$, if $(X[\leq])^{2}$ is normal then $X$ must be a Q-set. This would vacuously follow from a positive answer to the following possibly more interesting question:
Question 1.
Is it relatively consistent with ZFC that for any uncountable set of reals $X$, $(X[\leq])^{2}$ is not normal?
By Corollary 2.1 if Baumgartner’s Axiom holds and there are no $Q$-sets, then $X([\leq])^{2}$ is never normal for any uncountable set of reals $X$. So we ask,
Question 2.
Does Baumgartner’s Axiom imply that there is a $Q$-set?

### v2

@@
-Normality in the square of the Sorgenfrey line
+Submitted toTopology Proceedings
 Abstract.
@@
 Przymusiński proved that if X is a Q-set then $(X[\leq])^{2}$ is normal.
-While the converse is not in general true we consider examples of sets of the reals for which $(X[\leq])^{2}$ is normal or just pseudo-normal.
-For example, if $X$ is a $\lambda$ set, then $(X[\leq])^{2}$ is pseudo-normal but assuming CH there is an $X$ concentrated on a countable dense subset (so not a $\lambda$-set) but still $(X[\leq])^{2}$ is normal.
+Todorčević proved if X is an entangled set, then all finite power of $X[\leq]$ are hereditarily Lindelöf, hence $(X[\leq])^{2}$ is normal.
+We construct from CH a subset X of reals, such that X is not a $\lambda$-set (hence not a Q-set), X is not 2-entangled, but still $(X[\leq])^{2}$ is normal.
+We will also prove the natural analogue of Przymusiński’s theorem for $\lambda$-sets, namely, if X is a $\lambda$-set then $(X[\leq])^{2}$ is pseudo-normal.
 Key words and phrases:
-$Q$-set, $\lambda$-set, Sorgenfrey topology
+$Q$-set, $\lambda$-set, entangled sets, Sorgenfrey topology, normality
 2020 Mathematics Subject Classification:
-Primary:
+Primary: 54D15, 54G20 Secondary: 03E50, 03E05, 03E75
+The first author acknowledges support from NSERC Grant RGPIN-2025-06324.
 1. Introduction
-A $Q$-set is a subset of the reals with the property that every subset is $G_{\delta}$ (equivalently $F_{\sigma}$) in the relative subspace topology.
-The notion was first introduced in [hausdorff1933] and $Q$-sets were used to produce consistent counterexamples to the Normal Moore Space Conjecture [Bing1951].
-For example, if $X\subseteq 2^{\omega}$, the set of branches $b_{x}=\{x\upharpoonright n\}$ in $2^{<\omega}$ is an almost disjoint family and the corresponding $\Psi$-space on $2^{<\omega}$ is normal if and only if $X$ is a $Q$-set.
-Similarly, if $X$ is a subset of the $x$-axis in the upper half plan of ${\mathbb{R}}^{2}$, then the Moore-Niemitzky Plane topology (aka the bubble space) is normal if and only if $X$ is a Q-set.
-For more on these examples see [T].
-On the other hand Prymusiński [T.Przymusinski1973] proved that if $X$ is a Q-set endowed with the Sorgenfrey Topology then $X^{2}$ is normal.
+A $Q$-set is an uncountable subset of the reals with the property that every subset is $G_{\delta}$ (equivalently, $F_{\sigma}$) in the relative Euclidean subspace topology.
+The notion was first introduced in [4] and much later $Q$-sets were used to produce consistent counterexamples to the Normal Moore Space Conjecture [2].
+For example, if $X\subseteq 2^{\omega}$, the set of branches $b_{x}=\{x\upharpoonright n:n<\omega\}$ in $2^{<\omega}$ is an almost disjoint family, and the corresponding $\Psi$-space on $2^{<\omega}$ is normal if and only if $X$ is a $Q$-set.
+Similarly, if $X$ is a subset of the $x$-axis in the upper half plane of ${\mathbb{R}}^{2}$, then the Moore-Niemyitzki Plane topology (a.k.a.
+the bubble space) is normal if and only if $X$ is a Q-set.
+For more on these examples see [7].
+On the other hand Prymusiński [6] proved that if $X$ is a Q-set endowed with the Sorgenfrey Topology then $X^{2}$ is normal.
 Recall, the Sorgenfrey line is the real line endowed with the topology, also known as the lower limit topology, generated by using the half-open intervals $[a,b)$ as a basis.
 For a set of reals $X$ we will denote $X[\leq]$ as the resulting topology taking $X$ as a subspace of the Sorgenfrey line.
-We first observe that the converse to Przymusiński’s theorem does not hold: Todorčević observed that if $X$ is an entangled set of reals and is endowed with the Sorgenfrey topology, then every finite power is both hereditarily Lindelöf and hereditarily separable [Todorcevic1989].
-Recall that CH implies both the existence of entangled sets of reals and the non-existence of $Q$-set.
-There is the weaker notion of a set of reals $X$ being a $\lambda$-set.
-$X$ is a $\lambda$-set if every countable subset is $G_{\delta}$.
-While the existence of a $Q$-set of reals is independent, there are $\lambda$-sets in ZFC.
-Analagous to the $Q$-set results mentioned above, the $\Psi$-space over the set of branches of a subset $X\subseteq 2^{\omega}$ is pseudo-normal (meaning that every pair of closed sets, one of which is countable, can be separated by disjoint open sets) if and only if $X$ is a $\lambda$-set and also $X$ is a $\lambda$-set if and only if the Moore-Niemytzki plane over $X$ is pseudo-normal (see [vD])
-We provide some partial converses to Przymusiński’s theorem that underscore what is needed to prove establish normality of $(X[\leq])^{2}$.
+We first observe that the converse to Przymusiński’s theorem does not hold: Todorčević observed that if $X$ is an entangled set [8] of reals and is endowed with the Sorgenfrey topology, then every finite power is both hereditarily Lindelöf and hereditarily separable [9].
+Recall that CH implies both the existence of entangled sets of reals and the non-existence of $Q$-sets.
+So $(X[\leq])^{2}$ may be normal even if $X$ is not a $Q$-set.
+There is the weaker notion of a set of reals $X$ being a $\lambda$-set [5].
+$X$ is a $\lambda$-set if every countable subset is relatively $G_{\delta}$ in the Euclidean topology.
+While the existence of a $Q$-set of reals is independent of ZFC, there are $\lambda$-sets in ZFC.
+Analogous to the $Q$-set results mentioned above, the $\Psi$-space over the set of branches of a subset $X\subseteq 2^{\omega}$ is pseudo-normal (meaning that every pair of closed sets, one of which is countable, can be separated by disjoint open sets) if and only if $X$ is a $\lambda$-set and also $X$ is a $\lambda$-set if and only if the Moore-Niemytzki plane over $X$ is pseudo-normal (see [10]).
+We provide some partial converses to Przymusiński’s theorem underscoring necessary conditions for normality of $(X[\leq])^{2}$.
 We show that the expected analogous result for the Sorgenfrey topology on a $\lambda$-set holds.
-Namely, if $X$ is a $\lambda$-set of reals, then $(X[\leq])^{2}$ is pseudonormal.
-We also construct, assuming CH, a set of reals for which $X([\leq])^{2}$ is normal such that $X$ is concentrated on a countable dense subset, and so not a $\lambda$-set.
+Namely, if $X$ is a $\lambda$-set of reals, then $(X[\leq])^{2}$ is pseudo-normal.
+We also construct, assuming CH, a set of reals for which $X$ is concentrated on a countable dense subset (so not a $\lambda$-set) and so that $X$ is also not 2-entangled, but still $X[\leq]^{2}$ is normal.
 We will use the following terminology and notation throughout.
@@
 Let $X$ be a topological space.
-Let $A,B$ be subsets in $X$, AA has the covering property with respect to $B$ iff there exists a countable cover $\{U_{n}\}_{n<\omega}$ such that $\overline{U_{n}}\cap B=\emptyset$.
+Let $A,B$ be subsets in $X$, AA has the covering property with respect to $B$ iff there exists a countable open cover $\{U_{n}\}_{n<\omega}$ such that $\overline{U_{n}}\cap B=\emptyset$.
 If there is no ambiguity in the context, we just say $A$ has the covering property.
@@
 Let $z=(x,y)\in\mathbb{R}^{2}$, $r>0$ in $\mathbb{R}$, denote the basic Sorgenfrey neighbourhood of radius $r$ around $z$ to be $U(z,r)=[x,x+r)\times[y,y+r)$.
+Since we will have two topologies on $\mathbb{R}$, we will use the following notation for closure and compliments.
+Definition 1.3.
+For $X\subseteq\mathbb{R}$ be a subset of reals, we denote $\overline{X}^{E}$ as the Euclidean closure and $\operatorname{int}_{E}(X)$ as Euclidean interior.
+We denote $\overline{X}^{\sigma}$ and $\operatorname{int}_{\sigma}(X)$ to denote the Sorgenfrey closure and Sorgenfrey interior respectively.
+The Sorgenfrey topology is finer than the Euclidean topology, i.e. every Euclidean open set is also Sorgenfrey open but not vice versa.
 Definition 1.4.
 Let $A,B\subseteq\mathbb{R}$, $f:A\to B$, denote the graph of $f$ in $\mathbb{R}^{2}$ to be $G_{f}=\{(x,f(x)):x\in A\}$.
-All of our topological terminology is standard and we refer the reader to Engelking for any undefined notions.
+All of our topological terminology is standard and we refer the reader to Engelking [3] for any undefined notions.
 2. Q-sets and the Sorgenfrey square
@@
 Let $A,B$ be subsets of reals.
-For any strictly decreasing function $f:A\to B$, there is a countable $A_{0}$ such that $f\upharpoonright(A\setminus A_{0}):A\setminus A_{0}\rightarrow f(A\setminus A_{0})$ is a homeomorphism.
+For any monotone decreasing function $f:A\to B$, there is a countable set $B_{0}\subseteq B$ such that $f\upharpoonright(A\setminus f^{-1}(B_{0})):A\setminus f^{-1}(B_{0})\to f(A)\setminus B_{0}$ is a Euclidean homeomorphism.
+If, further, $f$ is strictly decreasing, then there is a countable $A_{0}\subseteq A$ such that $f\upharpoonright(A\setminus A_{0}):A\setminus A_{0}\rightarrow f(A\setminus A_{0})$ is a Euclidean homeomorphism.
 Lemma 2.2.
-Suppose $X$ is a set of reals. Suppose further that
+Suppose $X,Y$ are sets of reals. Suppose that
   1. (1)
-    there exists a strictly decreasing homeomorphism $f:X\to X$, and
+    there exists a strictly decreasing Euclidean homeomorphism $f:X\to Y$.
   2. (2)
-    letting $A\subseteq X$, define
+    For $A\subseteq X$, define
     $B$ $=X\setminus A$
@@
     $G_{f|_{B}}$ $=\{(b,f(b):b\in B)\}.$
-If there exists $\mathcal{U},\mathcal{V}\subseteq X^{2}$ Sorgenfrey open such that $\mathcal{U}\supseteq G_{f|_{A}},\mathcal{V}\supseteq G_{f_{B}}$ and $\mathcal{U}\cap\mathcal{V}=\emptyset$, then $A$ is relatively $F_{\sigma}$ in $X$.
+If there exists $\mathcal{U},\mathcal{V}\subseteq X\times Y$ Sorgenfrey open such that $\mathcal{U}\supseteq G_{f|_{A}},\mathcal{V}\supseteq G_{f_{B}}$ and $\mathcal{U}\cap\mathcal{V}=\emptyset$, then $A$ is relatively $F_{\sigma}$ in $X$ in the Euclidean topology.
 Proof.
 By assumption, let $\mathcal{U},\mathcal{V}$ be Sorgenfrey open sets such that $\mathcal{U}\supseteq G_{f|_{A}},\mathcal{V}\supseteq G_{f|_{B}}$.
-Note that $A=\cup_{n<\omega}A_{n}$ where $A_{n}=\{a\in A:U((a,f(a)),1/n)\subseteq\mathcal{U}\}$, we claim that
+Note that $A=\bigcup_{m<\omega}A_{m}$ where $A_{m}=\{a\in A:U((a,f(a)),1/m)\subseteq\mathcal{U}\}$, we claim that
 Claim 2.3.
-$\overline{A_{n}}^{E}\cap X\subseteq A$
-Proof.
-Assume otherwise, if $a_{n}\to b$ for $a_{n}\in A$ and $b\in B$.
+$\overline{A_{m}}^{E}\cap X\subseteq A$
+With this claim, we have $A=\bigcup_{m<\omega}(\overline{A_{m}}^{E}\cap X)$ which is a relatively Euclidean $F_{\sigma}$ set. So it suffices to prove the claim.
+Proof of 2.3.
+Assume otherwise, if $a_{n}\to b$ for $a_{n}\in A_{m}$ and $b\in B$.
 There exists an $r>0$ such that $U((b,f(b)),r)\subseteq\mathcal{V}$.
 By continuity of $f$, we have $f(a_{n})\to f(b)$ as well.
-Let $\epsilon=\min(1/n,r)$, Wlog there exists $a_{n_{0}}$ such that $b<a_{n_{0}}<b+\epsilon$ and $f(b)<f(a_{n_{0}})<f(b)+\epsilon$.
-Since $\epsilon<1/n$, $(a_{n_{0}},f(a_{n_{0}})+\epsilon)$ is a point in $U((a_{n_{0}},f(a_{n_{0}})),1/n)\subseteq\mathcal{U}$ However, we have
-$f(b)<f(a_{n_{0}})+\epsilon<f(b)+2\epsilon<f(b)+r$
-So $(a_{n_{0}},f(a_{n_{0}}))$ is also a point in $U((b,f(b)),r)\subseteq\mathcal{V}$, but $\mathcal{U}$ and $\mathcal{V}$ should have empty intersection hence this is a contradiction. ∎
-Now with this claim, we have $A=\cup_{n<\omega}(\overline{A_{n}}^{E}\cap X)$ which is a relatively $F_{\sigma}$ set. ∎
+Let $\epsilon<\min(1/m,r)$, without loss of generality there exists $a_{n_{0}}$ such that $b>a_{n_{0}}>b-\epsilon$ and $f(b)<f(a_{n_{0}})<f(b)+\epsilon$.
+Since $\epsilon<\min(1/m,r)$, we have
+$f(b)<\mathbf{f(a_{n_{0}})}<f(b)+\epsilon<f(b)+r$ $\implies\mathbf{f(a_{n_{0}})}\in[f(b),f(b)+r)$
+$a_{n_{0}}<\mathbf{b}<a_{n_{0}}+\epsilon<a_{n_{0}}+1/m$ $\implies\mathbf{b}\in[a_{n_{0}},a_{n_{0}}+1/m)$
+Where the first expression implies $(b,f(a_{n_{0}}))\in U((b,f(b)),r))\subseteq\mathcal{V}$ However, the second expression implies $(a_{n_{0}},f(a_{n_{0}}))$ is also a point in $U((a_{n_{0}},f(a_{n_{0}})),1/m)\subseteq\mathcal{U}$, but $\mathcal{U}$ and $\mathcal{V}$ should have empty intersection hence this is a contradiction.
+∎
 We now prove the following partial converse to Przymusiński’s theorem:
 Proposition 2.4.
-Let $(X[\leq])^{2}$ be normal. If there exists a strictly decreasing homeomorphism $f:X\to X$, then $X$ is a $Q$-set.
+Let $(X[\leq])^{2}$ be normal. If there exists a strictly decreasing Euclidean homeomorphism $f:X\to X$, then $X$ is a $Q$-set.
 Proof.
-Since $f$ is a homeomorphism, its graph $G_{f}$ is closed and discrete.
-Hence any two subsets of $G_{f}$ is closed.
-In particular, take any $A\subseteq X$, $G_{f|_{A}}$ and $G_{f|_{X\setminus A}}$ can be separated by open sets.
-By 2.2, $A$ is relatively $F_{\sigma}$ in $X$.
+Since $f$ is an Euclidean homeomorphism, its graph $G_{f}$ is Euclidean closed and so Sorgenfrey closed.
+Since $f$ is strictly decreasing, $G_{f}$ is also Sorgenfrey discrete.
+Hence any subset of $G_{f}$ is Sorgenfrey closed.
+In particular, for any $A\subseteq X$, $G_{f|_{A}}$ and $G_{f|_{X\setminus A}}$ can be separated by Sorgenfrey open sets (by our normality assumption).
+By 2.2, $A$ is relatively Euclidean $F_{\sigma}$ in $X$.
 Hence $X$ is a $Q$-set.
@@
 Proof.
-By 2.1, there exists a countable set $C$ such that $f|_{X\setminus C}:X\setminus C\to X\setminus C$ is a homeomorphism.
+By 2.1, there exists a countable set $C$ such that $f|_{X\setminus C}:X\setminus C\to f(X\setminus C)$ is a Euclidean homeomorphism.
 Then let $g=f|_{X\setminus C}$, the graph $G_{g}$ of $g$ is closed and discrete.
@@
 In particular, let $X^{\prime}=X\setminus C$, take any subset $A\subseteq X^{\prime}$, $G_{g|_{A}}$ and $G_{g|_{X\prime\setminus A}}$ can be separated by open sets $\mathcal{U},\mathcal{V}$.
-Take their intersection to $(X^{\prime})^{2}$, $\mathcal{U}\cup(X^{\prime})^{2}$ and $\mathcal{V}\cup(X^{\prime})^{2}$ are still Sorgenfrey open in $(X^{\prime})^{2}$.
-By 2.2, $A$ is relatively $F_{\sigma}$ in $X^{\prime}$, hence $X^{\prime}$ is a $Q$ set, and it only lacks countably many points.
+Take their intersection to $(X^{\prime})^{2}$, $\mathcal{U}\cap(X^{\prime})^{2}$ and $\mathcal{V}\cap(X^{\prime})^{2}$ are still Sorgenfrey open in $(X^{\prime})^{2}$.
+By 2.2, $A$ is relatively Euclidean $F_{\sigma}$ in $X^{\prime}$, hence $X^{\prime}$ is a $Q$ set as required.
 ∎
-Baumgartner proved that PFA implies that any two $\aleph_{1}$-dense sets of reals are order isomorphic [Baumgartner1973].
+Baumgartner proved that PFA implies that any two $\aleph_{1}$-dense sets of reals are order isomorphic [1].
 This consequence of PFA has come to be referred to as Baumgartner’s Axiom (BA).
@@
 Proof.
-Any uncountable $X$ contains an $\aleph_{1}$-dense subset.
+Any uncountable $X$ contains an $\aleph_{1}$-dense subset $X^{\prime}$.
 By BA it is order isomorphic to its reverse order.
-So by the proof of Proposition 2.5 there is a countable $C$ such that $X\setminus C$ must be a $Q$-set.∎
+So by the proof of Proposition 2.5 there is a countable $C$ such that $X^{\prime}\setminus C$ must be a $Q$-set.∎
 For the remainder of this section we focus on constructing, from CH, an uncountable set of reals whose square, under the Sorgenfrey topology, is normal.
@@
 Lemma 2.7.
-Let $f:A\to\mathbb{R}$ be strictly decreasing on uncountable $A\subseteq\mathbb{R}$, then $f$ can be extended to strictly decreasing function on $\overline{A}\setminus C$ for some countable set $C$.
+Let $f:A\to\mathbb{R}$ be strictly decreasing on uncountable $A\subseteq\mathbb{R}$, then there is a strictly decreasing homeomorphism $F$ on $\overline{A}^{E}\setminus C$ for some countable set $C$ such that $f(a)=F(a)$ for all $a\in A\setminus C$.
 Proof.
-$f$ can be extended to a decreasing function on $\overline{A}$.
+$f$ can be extended to a decreasing function on $\overline{A}^{E}$.
 Note that this function is not necessarily strictly decreasing.
-By 2.1, we can remove a countable set $B_{0}$ such that $f$ is homeomorphism on $\overline{A}\setminus f^{-1}(B_{0})$ hence strictly decreasing.
+By 2.1, we can remove a countable set $B_{0}$ such that $f$ is a Euclidean homeomorphism $F$ on $\overline{A}^{E}\setminus f^{-1}(B_{0})$ hence strictly decreasing.
 We will show $f^{-1}(B_{0})$ is countable by showing for every $b\in B_{0}$, $f^{-1}(b)$ has at most 3 points.
@@
 but $f$ is strictly decreasing on $A$, so there can only be at most 1 point in $A$.
-Hence, there are no points in $\overline{A}\setminus A$.
+Hence, there are no points in $\overline{A}^{E}\setminus A$.
 But $y,z$ can not be both in $A$, we have a contradiction.
@@
 ∎
-Recall a set $L\subseteq\mathbb{R}$ is $\mathbf{n}$-entangled iff for any set of disjoint $n$-tuples $A\subseteq L^{n}$ of size $\omega_{1}$ and for every $s\in 2^{n}$ there exists $x,y\in A$ such that $\forall i\leq n(x_{i}<y_{i})\iff s_{i}=0$.
+We also recall definition of entangled sets from [8].
+Let $L\subseteq\mathbb{R}$, denote $(L)^{n}\subseteq L^{n}$ as the set of all $\langle x_{0},\dots,x_{n-1}\rangle\in L^{n}$, such that $x_{0}<x_{1}<\dots<x_{n}$.
+A set $L\subseteq\mathbb{R}$ is $\mathbf{n}$-entangled iff for any set of disjoint $n$-tuples $A\subseteq(L)^{n}$ of size $\omega_{1}$ and for every $s\in 2^{n}$ there exists $x,y\in A$ such that $\forall i\leq n,(x_{i}<y_{i})\iff s_{i}=0$.
 $L$ is entangled iff $L$ is $\mathbf{n}$-entangled for all $n<\omega$.
-Let $s\in 2^{n}$, $A\subseteq L^{n}$ is said to have type $s$ iff $\forall i\leq n(x_{i}<y_{i})\iff s_{i}=0$.
-The following lemma is due to [Todorcevic1985].
+Let $s\in 2^{n}$, $A\subseteq(L)^{n}$ is said to have type $s$ iff $\forall i\leq n(x_{i}<y_{i})\iff s_{i}=0$.
+The following lemma is proven in [8] (see property (5) in the proof of Theorem 1 in [8]).
 Lemma 2.8.
-[Todorcevic1985] Let $E\subseteq\mathbb{R}$ to have the following property:
-  • No one-to-one continuous function can be defined from $A\to E$ for any uncountable $A\subseteq E^{n}$, for any $n<\omega$
+[8] Let $E\subseteq\mathbb{R}$ be such that for any continuous function $f$ from a $G_{\delta}$ subset of ${\mathbb{R}}^{n}$ into $\mathbb{R}$}, $(n<\omega)$ there is an $\alpha<2^{\omega}$ such that
+$\forall\beta\geq\alpha,f(\{r_{\gamma}:\gamma<\beta\}^{n})\cap E\subseteq\{r_{\gamma}:\gamma<\beta\}$
 Then $E$ is entangled.
-Proof.
-Assume $E$ is not entangled, let $m<\omega$ be the least such that $E$ is $m$-entangled, but not $m+1$-entangled.
-Then there exists an uncountable $A\subseteq E^{m+1}$ such that $A$ is missing one type $s\in 2^{n}$, with $s=(s_{1},s_{2},\dots,s_{m},s_{m+1})$.
-Wlog, let $s_{m+1}=0$.
-Define $f:E^{m}\to E$ by $f(z)=r$ iff $z\text{\textasciicircum}r\in A$.
-We will show $f$ is continuous on some uncountable subset of $dom(f)$, which is not possible by the property $E$ has.
-For any $p\in dom(f)$, define
-$\omega(p)=\cap\{\overline{f(dom(f)\cap I)}:I\text{ open in }E^{m},p\in I\}$
-$\omega(p)$ is all the possible limit at the point $p$, if $|\omega(p)|=0$ or $1$, $f$ is continuous at $p$.
-Let $B=\{p\in dom(f):|\omega(p)|\geq 2\}$, assume $B$ is uncountable.
-By countable chain condition and symmetry, there exists an uncountable subset $B_{0}$ such that
-$\forall z\in B_{0},\exists r\in\omega(z)\;f(z)<d<r$
-Note that $B_{0}\subseteq E^{m}$, and since $E$ is $m$-entangled, there exists $u,v\in B_{0}$, $u=(u_{1},\dots,u_{m}),v=(v_{1},\dots,v_{m})$, such that $\forall i\leq m(u_{i}<v_{i})\iff s_{i}=0$. Let
-$I=\{z\in E^{m}:\forall i\leq m(u_{i}<z_{i})\iff s_{i}=0\}$
-II is an open set.
-However for all $z\in dom(f)\cap I$, since $s_{m+1}=0$ and $A$ is missing this type, we can only have $f(z)<f(u)<d$, and $\omega(v)\subseteq\overline{f(dom(f)\cap I)}\implies\omega(v)<d$ which is a contradiction.
-∎
 Theorem 2.9.
@@
   1. (1)
-    No one-to-one continuous function can be defined from $A\to E$ for any uncountable $A\subseteq E^{n}$, for any $n<\omega$.
+    For any continuous function $f$ from a $G_{\delta}$ subset of ${\mathbb{R}}^{n}$ into $E$, $(n<\omega)$ there is an $\alpha<2^{\omega}$ such that
+    $\forall\beta\geq\alpha,f(\{r_{\gamma}:\gamma<\beta\}^{n})\cap E\subseteq\{r_{\gamma}:\gamma<\beta\}$
   2. (2)
-    $E$ contains $\mathbb{Q}$ and $\mathbb{Q}$ is not a relative $G_{\delta}$-set in $E$.
-Moreover, $E$ can be chosen so that it is concentrated on $\mathbb{Q}$.
-I.e., $E\setminus U$ is countable for every open set $U\supseteq\mathbb{Q}$.
+    $E$ contains $\mathbb{Q}$ and is concentrated on $\mathbb{Q}$}.
+Recall that $E$ is concentrated on $\mathbb{Q}$ means that $E\setminus U$ is countable for every open set $U\supseteq\mathbb{Q}$, so $E$ is not a $\lambda$-set since $\mathbb{Q}$ is not a relative $G_{\delta}$-set in $E$.
 Proof.
 Define
-$\mathcal{F}=\{$ $f:U\to\mathbb{R}:f\text{ one-to-one continuous function on Borel set }U\subseteq\mathbb{R}^{n},n<\omega,$
-$\text{and let }U=\Pi_{i=1}^{n}U_{i},\text{we have }\cup_{i=1}^{n}U_{i}\text{ uncountable}\}$
+$\mathcal{F}=\{$ $f:U\to\mathbb{R}:U\subseteq\mathbb{R}^{n},n<\omega,U\text{ is }G_{\delta}\}\}$
 $\mathcal{A}=\{$ $A\subseteq\mathbb{R}:A\text{ is open},\ A\supseteq\mathbb{Q}\}$
@@
 $\mathcal{F}=\{f_{\alpha}:\alpha<\omega_{1}\},\mathcal{A}=\{A_{\alpha}:\alpha<\omega_{1}\}.$
-For every $\alpha<\omega_{1}$ define the following set:
-$C_{\alpha}$ $=\bigcap_{\beta\leq\alpha}A_{\alpha}\setminus\mathbb{Q}.$
-Note that $C_{\alpha}$ is uncountable by the Baire Category Theorem. We now define elements $x_{\alpha}\in\mathbb{R}$ recursively as follows:
-Stage 0: Take $x_{0}\in C_{0}$ such that $f_{0}(\{x_{0}\}^{n})\not=x_{0}$
-Stage $\bm{\alpha<\omega_{1}}$: We have $x_{\beta}$ defined already for all $\beta<\alpha$, define
-$X_{\alpha}$ $=\{x_{\beta}:\beta<\alpha\}$
-$I_{\alpha}$ $=\{f_{\beta}(x_{\gamma_{1}},x_{\gamma_{2}},\dots,x_{\gamma_{n}}):\beta\leq\alpha,\gamma_{i}<\alpha,i\leq n\}$
-$P_{\alpha}$ $=\cup\{f_{\beta}^{-1}(x_{\gamma}):\beta\leq\alpha,\gamma<\alpha\}$
-Note that all of the above sets are countable, and $C_{\alpha}$ is uncountable.
-Hence, we can choose $x_{\alpha}\in C_{\alpha}\setminus(X_{\alpha}\cup I_{\alpha}\cup P_{\alpha})$, such that either $x_{\alpha}\not\in\cup dom(f_{\alpha})$ or $f_{\alpha}([X_{\alpha}]^{n}\cap dom(f))\not\ni x_{\alpha}$.
+For every $\alpha<\omega_{1}$ let $n_{\alpha}$ be the parity of the function $f_{\alpha}$ and define the following set:
+$C_{\alpha}$ $=\bigcap_{\beta\leq\alpha}A_{\beta}\setminus\mathbb{Q}.$
+Note that $C_{\alpha}$ is uncountable by the Baire Category Theorem. We now define elements $x_{\alpha}\in\mathbb{R}$ and countable sets $I_{\alpha}\subseteq{\mathbb{R}}$ recursively as follows:
+Stage 0: Let $I_{0}=f_{0}({\mathbb{Q}}^{n_{0}})$ and take $x_{0}\in C_{0}\setminus I_{0}$.
+Stage $\bm{\alpha<\omega_{1}}$: We have $x_{\beta}$ and $I_{\beta}$ defined already for all $\beta<\alpha$. For any $\beta\leq\omega_{1}$, let $n_{\beta}$ be the parity of $f_{\beta}$, and let
+$X_{\alpha}$ $={\mathbb{Q}}\cup\{x_{\beta}:\beta<\alpha\}$
+$I_{\alpha}$ $=\bigcup_{\beta\leq\alpha}f_{\beta}(X_{\alpha}^{n_{\beta}})$
+Note that all of the above sets are countable, and $C_{\alpha}$ is uncountable. Hence, we can choose $x_{\alpha}\in C_{\alpha}\setminus(X_{\alpha}\cup I_{\alpha})$.
 Finally we define $E=\mathbb{Q}\cup\{x_{\alpha}:\alpha<\omega_{1}\}$.
-Verify Property 1: Fix $n<\omega$, for any one to one continuous function $g:A\to E$ defined on some uncountable set $A\subseteq E^{n}$, $g$ can always be extended to a continuous function on a Borel set.
-, i.e., there is an $\alpha_{0}$ such that $f_{\alpha_{0}}\upharpoonright A=g$.
-However, by choice of $x_{\alpha}$, for all $\alpha>\alpha_{0}$, we have that $f_{\alpha_{0}}(\{x_{\gamma}:\gamma<\alpha\}^{n})$ can not take on any values in $\{x_{\beta}:\beta>\alpha_{0}\}$.
-Indeed for all $\beta>\alpha_{0}$, if $\alpha_{0}<\beta<\alpha$, we have $x_{\alpha}\not\in P_{\alpha}\implies x_{\alpha}\not\in\cup f^{-1}(x_{\beta})\implies f_{\alpha_{0}}(\{x_{\gamma}:\gamma<\alpha\}^{n})\neq x_{\beta}$.
-If $\alpha_{0}<\alpha<\beta$, we have $x_{\beta}\not\in I_{\beta}\implies x_{\beta}\not\in f(\{x_{\gamma}:\gamma<\beta\}^{n})\implies x_{\beta}\not\in f(\{x_{\gamma}:\gamma<\alpha\}^{n})$.
-So for all $z\in E^{n}$ such that $z_{i}=x_{\alpha}$ for some $i\leq n$, $f_{\alpha_{0}}(z)\in\mathbb{Q}\cup\{x_{\beta}:\beta\leq\alpha_{0}\}$ which is a countable set.
-Hence $g^{-1}(E)$ is countable contradicting the fact that $g$ was defined on the uncountable set $A$.
-Verify Property 2: That $X$ is concentrated on $\mathbb{Q}$ follows easily from the fact that for every $\alpha$, $X\setminus A_{\alpha}\subseteq\{x_{\beta}:\beta<\alpha\}$.
-In particular, $\mathbb{Q}$ is not $G_{\delta}$ in $X$ and so $X$ is not a $\lambda$-set.
-By 2.7, the set $E$ we constructed in 2.8 is entangled and in particular $2$-entangled. By 2.6 this implies no monotone functions can be defined on $E$. Hence we have the following result.
+Verify Property (1): Fix $n<\omega$, for any continuous function $g:A\to E$, by choice of $x_{\alpha}$, we have that for any $\beta\geq\alpha$, $f_{\alpha_{0}}(\{x_{\gamma}:\gamma<\alpha\}^{n})$ can not take on any values in $\{x_{\beta}:\beta\geq\alpha_{0}\}$.
+Indeed for all $\beta>\alpha_{0},$ $x_{\beta}\not\in I_{\alpha}\supseteq f_{\alpha_{0}}((\{x_{\alpha}:\alpha<\alpha_{0}\}\cup\mathbb{Q})^{n})$
+Verify Property (2): That $E$ is concentrated on $\mathbb{Q}$ follows easily from the fact that for every for every open set containing $\mathbb{Q}$} is enumerated as $A_{\alpha}$ and so $C_{\beta}\subseteq A_{\alpha}$ for all $\beta\geq\alpha$.
+As mentioned above, the set $E$ constructed is not a $\lambda$-set so also not a $Q$-set.
+By 2.8, the set $E$ is also entangled and in particular $2$-entangled. By 2.7 this implies no monotone functions can be defined on an uncountable subset of $E$.
+For our next observation, we will use the following result, proven by Przymusiński (see the proof of Theorem 1 in [6]).
+Lemma 2.10.
+Let $X\subseteq\mathbb{R}$ be any subset of reals.
+Fix any Sorgenfrey closed set $A,B\subseteq X^{2}$, and a Sorgenfrey open cover $\mathcal{A}$ of $A$.
+For any $z\in B^{c}$ choose $n_{z}\in\mathbb{N}$ such that $U(z,1/n_{z})\cap B=\emptyset$, this is possible since $B$ is Sorgenfrey closed.
+Let $A\cap\overline{B}^{E}=\bigcup_{q\in\mathbb{Q}}A_{q}$ where
+$A_{q}=\{z=(x,y)\in A_{1}:x+y<q<x+y+1/n_{z}\}$
+For any $q\in\mathbb{Q}$, we have
+  1. (1)
+    $A_{q}=G_{q}\cup L_{q}$, where $L_{q}$ can be covered by countably many Sorgenfrey open sets of $\mathcal{A}$ and $G_{q}$ is a graph of some strictly decreasing function,
+  2. (2)
+    If every $G_{q}$ can be covered by countably many Sorgenfrey open sets in $\mathcal{A}$, then $X[\leq]^{2}$ is normal.
+Recall, Todorčević noted [9] that if $E$ is entangled, then $E[\leq]$ is hereditarily Lindelöf in all finite powers. In relation to this we have:
 Theorem 2.11.
-Let $E$ be any subset of $\mathbb{R}$} that admits no uncountable strictly monotone functions (e.g., the set constructed in Theorem 2.8), then $(E[\leq])^{2}$ is normal.
+Let $E$ be any subset of $\mathbb{R}$} that admits no uncountable strictly decreasing functions (e.g., the set constructed in Theorem 2.9), then $(E[\leq])^{2}$ is normal.
 Proof.
-Let $A,B$ be closed in $(E[\leq])^{2}$.
-Recall the shoelace lemma: to separate $A$ and $B$ it suffices to cover each set by a countable family of open sets whose closures are disjoint from the other closed set.
-By symmetry, it suffices to find a countable cover $\{U_{n}\}_{n<\omega}$ of $A$ such that $\overline{U_{n}}\cap B=\emptyset$.
-Recall we defined $A^{\prime}\subseteq A$ to have the covering property (with respect to $B$) if there is a countable open cover $\{U_{n}\}$ such that $\overline{U_{n}}\cap B=\emptyset$
-Clearly, we have:
-Lemma 2.10.
-If $A^{\prime}\subseteq A$ is Lindelof, then $A^{\prime}$ has the covering property.
-For all $z\in B^{c}$ choose $n_{z}\in\mathbb{N}$ such that $U(z,1/n_{z})\cap B=\emptyset$, this is possible since $B$ is Sorgenfrey closed.
-By letting $A_{1}=A\cap\overline{B}^{E}$ and $A_{2}=A\setminus\overline{B}^{E}$, we have
-Observation 1.
-$A=A_{1}\cup A_{2}$
-Claim 2.11.
-$A_{2}$ has the covering property.
-Proof.
-Note that $A_{2}\subseteq E\setminus cl_{E}(B)=U$ and the latter is Euclidean open.
-Since $E^{2}$ is hereditarily Lindelöf with respect to the Euclidean topology, $U$ can be written as the countable union of basic open sets $B_{n}$ such that $\overline{B_{n}}\cap B=\emptyset$ for each $n$ so, since the Sorgenfrey topology is finer than the Euclidean, $A_{2}$ has the covering property.
+Following the notation of 2.10 (1): $G_{q}$ is the graph of a strictly decreasing function, but by hypothesis on $E$, each $G_{q}$ is countable, hence Lindelöf.
+By 2.10 (2), $(E[\leq])^{2}$ is normal.
 ∎
-Claim 2.12.
-$A_{1}$ has the covering property.
-Proof.
-We will break down $A_{1}$ into countably many pieces, each of which is the graph of a monotone decreasing function and prove every piece has the covering property. Define
-$A_{q}=\{z=(x,y)\in A_{1}:x+y<q<x+y+1/n_{z}\}$
-Denote $l_{q}=\{(x,y)\in X^{2}:x+y=q\}$, $A_{q}$ is just the collection of points $z$ which has the straight line $l_{q}$ passing through the interior of its nbhd $U(z,1/n_{z})$. We see that
-Observation 2.
-$A_{1}=\cup_{q\in\mathbb{Q}}A_{q}$
-So now it suffices to show that each $A_{q}$ has the covering property. First note that
-Claim 2.13.
-(non-increasing property) For all $z_{1}=(x_{1},y_{1}),z_{2}=(x_{2},y_{2})\in A_{q}$, if $x_{1}<x_{2}$, then $y_{2}\leq y_{1}$.
-Proof.
-For otherwise we have $x_{1}<x_{2}$ and $y_{1}<y_{2}$.
-Since the line $l_{q}$ passes through both nbhds $U(z_{1},1/n_{z_{1}}),U(z_{2},1/n_{z_{2}})$, this means $z_{2}$ is in the Euclidean interior of $U(z_{1},1/n_{z_{1}})$.
-However, $\operatorname{int}_{E}U(z_{1},1/n_{z_{1}})$ is a Euclidean open set, and since $z_{2}\in\overline{B}^{E}$, we have that $\operatorname{int}_{E}U(z_{1},1/n_{z_{1}})$ also contains a point from $B$, which is a contradiction, because we chose $n_{z_{1}}$ so that its nbhd $U(z_{1},1/n_{z_{1}})$ is disjoint from $B$.
-∎
-Claim 2.14.
-The sets $M_{1}=\{x\in p_{0}(A_{q}):|p_{0}^{-1}(x)|>1\},M_{2}=\{x\in p_{1}(A_{q}):|p_{1}^{-1}(x)|>1\}$ are countable.
-Proof.
-For each $x\in M_{1}$, fix $\alpha_{x}<\beta_{x}\in p_{0}^{-1}(x)$.
-And for any $x_{1}<x_{2}\in M_{1}$, we have $\beta_{x_{2}}\leq\alpha_{x_{1}}$ by the previous claim.
-And so the intervals $(\alpha_{x_{1}},\beta_{x_{1}})$ and $(\alpha_{x_{2}},\beta_{x_{2}})$ are disjoint.
-Hence, we have an uncountable set of pairwise disjoint open intervals on $\mathbb{R}$}, contradiction.
-The proof for $M_{2}$ is the same.
-∎
-Let $L_{1}=p_{0}^{-1}(M_{1}),L_{2}=p_{1}^{-1}(M_{2})$, and $L=L_{1}\cup L_{2}$, $K=A_{q}\setminus L$, we see that
-Observation 3.
-$A_{q}=K\cup L$
-Claim 2.15.
-$L$ has the covering property.
-Proof.
-We know $L_{1},L_{2}$ are Lindelof since by Claim 2.14 each is a countable union of Lindelof subspaces. And by 2.10, $L$ has the covering property. ∎
-Now it suffices to show $K$ has the covering property.
-Claim 2.16.
-$K$ has the covering property
-Proof.
-Note that $K$ is now a graph of some strictly decreasing function $g$. By the definition of $E$, such $g$ can only be defined on a countable set, hence the graph is countable hence Lindelof. ∎
-This shows that each $A_{q}$ has the covering property completing the proof that $A_{1}$ has the covering property. ∎
-Now with all the observations and claims, we obtain a countable cover of $A$ witnessing that $A$ has the covering property and so by the shoelace lemma, $A$ and $B$ can be separated by open sets, hence $E^{2}$ is normal.
-∎
-Now one may try to categorize all sets $E\subseteq\mathbb{R}$ such that $(E[\leq])^{2}$ into either $Q$-set or a $2$-entangled set. However, the following example shows that is not the case.
+Now one may ask if every $E\subseteq\mathbb{R}$ such that $(E[\leq])^{2}$ is normal is either a $Q$-set or is $2$-entangled. However, the following example shows that is not the case.
 Theorem 2.12.
@@
 Proof.
-Let $f:\mathbb{R}\to\mathbb{R}$ be any one-to-one increasing function such that $dom(f)$ is of cardinality $\aleph_{1}$ and $\{x\in dom(f):f(x)\neq x\}$ is of cardinality $\aleph_{1}$.
-Wlog, we can assume $f(x)\neq x$ for all $x\in dom(f)$.
-We will preserve this function $f$ while making it concentrated on rationals and killing all the strictly decreasing functions on uncountable sets.
-Let $G_{f}={(u_{\alpha},v_{\alpha}):\alpha<\omega_{1}}$.
-As in 2.8, let $\mathcal{F}=\{f_{\alpha}:\alpha<\omega_{1}\}$ be the set of strictly decreasing functions, $\mathcal{A}=\{A_{\alpha}:\alpha<\omega_{1}\}$ be the set of open sets containning $\mathbb{Q}$ and for each $\alpha$, $C_{\alpha}=\cap_{\beta\leq\alpha}A_{\alpha}\setminus\mathbb{Q}$.
+We will construct our set $E$ by preserving the strictly increasing function $x\mapsto x+1$ while at the same time making it concentrated on $\mathbb{Q}$} and killing all the strictly decreasing functions on uncountable sets.
+As in 2.9, let $\mathcal{F}=\{f_{\alpha}:\alpha<\omega_{1}\}$ enumerate all strictly decreasing functions defined on Borel subsets of the reals, $\mathcal{A}=\{A_{\alpha}:\alpha<\omega_{1}\}$ an enumeration of all Euclidean open sets containing $\mathbb{Q}$ and, for each $\alpha<\omega_{1}$, let $C_{\alpha}=\cap_{\beta\leq\alpha}A_{\alpha}\setminus\mathbb{Q}$.
 We now define elements $x_{\alpha}\in\mathbb{R}$ recursively as follows:
-Stage 0: Take $x_{0}\in C_{0}\setminus\{u_{0},v_{0}\}$ such that $f_{0}(x_{0})\not=x_{0}$
+Stage 0: Take $x_{0}\in C_{0}$ such that $f_{0}(x_{0})\not=x_{0}$ and $x_{0}\not\in f_{0}(\mathbb{Q})\cup f_{0}^{-1}(\mathbb{Q})$
 Stage $\bm{\alpha<\omega_{1}}$: We have $x_{\beta}$ defined already for all $\beta<\alpha$, define
-$X_{\alpha}$ $=\{x_{\beta}:\beta<\alpha\}\cup\{u_{\beta}:\beta<\alpha\}\cup\{v_{\beta}:\beta<\alpha\}$
+$X_{\alpha}$ $=\{x_{\beta}:\beta<\alpha\}\cup\{x_{\beta}+1:\beta<\alpha\}\cup\mathbb{Q}$
 $I_{\alpha}$ $=\{f_{\beta}(x):\beta\leq\alpha,x\in X_{\alpha}\}$
 $P_{\alpha}$ $=\{f_{\beta}^{-1}(x):\beta\leq\alpha,x\in X_{\alpha}\}$
-Note that all of the above sets are countable, and $C_{\alpha}$ is uncountable.
-Hence, we can choose $x_{\alpha}\in C_{\alpha}\setminus(X_{\alpha}\cup I_{\alpha}\cup P_{\alpha})$, such that either $x_{\alpha}\not\in dom(f_{\alpha})$ or $f_{\alpha}(X_{\alpha})\not\ni x_{\alpha}$.
-Finally we define $E=\mathbb{Q}\cup\{x_{\alpha}:\alpha<\omega_{1}\}\cup\{u_{\alpha}:\alpha<\omega_{1}\}\cup\{v_{\alpha}:\alpha<\omega_{1}\}$.
-Now $E$ is not entangled since the uncountable one-to-one increasing function $f$ exists. $(2)$ and $(3)$ follows exactly as in 2.8. ∎
+Note that all of the above sets are countable, and $C_{\alpha}$ is uncountable. Hence, we can choose $x_{\alpha}\in C_{\alpha}\setminus(X_{\alpha}\cup I_{\alpha}\cup P_{\alpha})$.
+Finally we define $E=\mathbb{Q}\cup\{x_{\alpha}:\alpha<\omega_{1}\}\cup\{x_{\alpha}+1:\alpha<\omega_{1}\}$.
+Verify Property (1): $E$ is not 2-entangled since the uncountable one-to-one increasing function $x\mapsto x+1$ exists.
+Verify Property (2): This follows exactly as in 2.9.
+Verify Property (3): We will show $(E[\leq])^{2}$ is normal using 2.11.
+Fix any strictly decreasing function $g$ defined on subsets of $E$.
+By 2.7, there exists $F$ a strictly decreasing homeomorphism defined from a subset of $E$ into $E$ such that $dom(F)\setminus dom(g)$ is countable.
+There exists an $\alpha_{0}<\omega_{1}$ such that $F=f_{\alpha_{0}}$ on $dom(F)$.
+We will show the $f_{\alpha_{0}}$ can only be defined on a countable set, hence $dom(g)$ is countable and $E$ admits no strictly decreasing function on uncountable subsets.
+Indeed, for all $\beta>\alpha_{0}$, $f_{\alpha_{0}}(x_{\beta})$ can not be mapped to $\{x_{\beta}:\beta\geq\alpha_{0}\}\}$.
+For all $\gamma\geq\alpha_{0}$,
+[…]
+$x_{\beta}\not\in P_{\beta}\ni f_{\alpha_{0}}^{-1}(x_{\gamma})\implies f_{\alpha_{0}}(x_{\beta})\neq x_{\gamma}$
+[…]
+$x_{\gamma}\not\in P_{\gamma}\ni f(x_{\beta})\implies x_{\gamma}\neq f_{\alpha_{0}}(x_{\beta})$
+For the same reason, $f_{\alpha_{0}}(x_{\beta})$ cannot be mapped to $\{x_{\beta}+1:\beta\geq a_{0}\}$. Since $f_{\alpha_{0}}$ is one-to-one, $f_{\alpha_{0}}$ has countable domain. ∎
 3. $\lambda$-sets and the Sorgenfrey square
 Theorem 3.1.
-If $X$ is a $\lambda$-set, then $(X[\leq])^{2}$ is pseudonormal.
+If $X$ is a $\lambda$-set, then $(X[\leq])^{2}$ is pseudo-normal.
 Proof.
 Let $A,B$ be closed sets and assume $B$ is countable.
-We will again use shoelace lemma, and so suffices to show that $A$ has the covering property with respect to $B$.
-Since $B$ is countable it clearly has the covering property with respect to $A$.
-We begin by defining the following sets:
+Recall the shoelace lemma: to separate $A$ and $B$ it suffices to cover each set by a countable family of open sets whose closures are disjoint from the other closed set.
+It suffices to show that $A$ has the covering property with respect to $B$. Since $B$ is countable it clearly has the covering property with respect to $A$. We begin by defining the following sets:
 $B_{0}$ $=p_{0}(B)$
@@
 $W$ $=A\setminus(A_{0}\cup A_{1})$
-Note that $B_{0},B_{1}$ are countable, so $A_{0}$ and $A_{1}$ are Lindelöf. Hence $A_{0}$ and $A_{1}$ have the covering property. Now it suffices to prove
+Note that $B_{0}$ is countable, then $p_{0}^{-1}(B_{0})=\bigcup_{b\in B_{0}}p_{0}^{-1}(b)$ is a countable union of Lindelöf spaces which is still Lindelöf.
+Since $A$ is Sorgenfrey closed, $A_{0}$ is a closed subspace of a Lindelöf space, which is still Lindelöf.
+By the similar reason, $A_{1}$ is Lindelöf as well.
+Hence both $A_{0}$ and $A_{1}$ have the covering property.
+Now it suffices to prove
 Claim 3.2.
@@
 Proof of Claim.
-We know that there are families $\{F_{n}:n\in\omega\}$ and $\{H_{n}:n\in\omega\}$ of relatively closed (in the Eucliden topology) subsets of $X$ such that
+We know that there are families $\{F_{n}:n\in\omega\}$ and $\{H_{n}:n\in\omega\}$ of relatively closed (in the Euclidean topology) subsets of $X$ such that
 $p_{0}(W)\subseteq p_{0}(B)^{c}$ $=\cup_{n\in\omega}F_{n}$
@@
 $W_{n,m}=\bigcup\{U(z,s_{z}):\ z=(x,y)\in W,x\in F_{n},y\in H_{m}\}.$
-We see $W\subseteq\cup_{n,m}W_{n,m}$ and each $W_{n,m}$ is Sorgenfrey open. We know that $W_{n,m}\cap B=\emptyset$, so it suffices to show $(\overline{W_{n,m}}\setminus W_{n,m})\cap B=\emptyset$.
+We see $W\subseteq\cup_{n,m}W_{n,m}$ and each $W_{n,m}$ is Sorgenfrey open.
+We know that $W_{n,m}\cap B=\emptyset$, so it suffices to show $(\overline{W_{n,m}}^{\sigma}\setminus W_{n,m})\cap B=\emptyset$.
 Assume otherwise.
@@
 Hence $\overline{W_{n,m}}\cap B=\emptyset$, we have the desired covering. ∎
-From the previous example, we know that it is possible for $(X[\leq])^{2}$ to be normal, while $X$ is not a $\lambda$-set.
-However, we do have the following partial converse analogous to Proposition 2.4.
+From the previous example, we know that $(X[\leq])^{2}$ may be normal, while $X$ is not a $\lambda$-set. However, we do have the following partial converse analogous to Proposition 2.4.
 Proposition 3.3.
@@
 Proof.
-Since $f$ is a homeomorphism, the graph $G_{f}$ is closed and discrete, any subset of $G_{f}$ is closed in $X^{2}$.
-Let $A$ be any countable subset, we have $G_{f|_{A}}$ is countable as well, hence $G_{f|_{A}}$ and $G_{f|_{X\setminus A}}$ are closed and can be separated by open sets.
+Since $f$ is strictly decreasing we have that the graph $G_{f}$ is closed and discrete in $(X[\leq])^{2}$.
+Therefore any subset of $G_{f}$ is closed.
+Let $A$ be any countable subset, we have $G_{f|_{A}}$ is countable as well, hence $G_{f|_{A}}$ and $G_{f|_{X\setminus A}}$ are closed and can be separated by Sorgenfrey open sets.
 Then by 2.2, $X\setminus A$ is relatively $F_{\sigma}$ in $X$, hence $X$ is a $\lambda$-set.
@@
 Question 1.
-Is it relatively consistent with ZFC that for any uncountable set of reals $X$, $(X[\leq])^{2}$ is not normal?
+Is it relatively consistent with ZFC that $(X[\leq])^{2}$ is never normal for $X$ an uncountable set of reals?
 By Corollary 2.6 if Baumgartner’s Axiom holds and there are no $Q$-sets, then $X([\leq])^{2}$ is never normal for any uncountable set of reals $X$. So we ask,